OpenStudy (anonymous):

f(x)=x^(3)+3x, find a. f(2+h)-f(2) b. f(1/4) c. f(√2) d f(1+h) e. f(1) f. f(1+h)-f(1)

OpenStudy (e.mccormick):

This is one where you substitute in and solve. Because you have the function defined, this should not be too hard. \[f(x)=x^3+3x\] Now plug (a.) into this: \[f(2+h)=(2+h)^3+3(2+h)\] and \[f(2)=(2)^3+3(2)\] therefore \[f(2+h)-f(2)= (2+h)^3+3(2+h)-[(2)^3+3(2)]\] now reduce that to simplest form. Do a simialr process for the rest.